# Annealing Approximation in Master-Node Network Model

### v1

@@ B Appendix – case $\overline{A^{(1)}}=0$ and $A^{(4)}>0$
Table 3: Division of the $(L,R)$ plane suitable for the construction of the flux diagram of the case $\overline{A^{(1)}}=0$ and $A^{(4)}>0$; the regions in Part III are defined in picture 8.
B.1 Part I (points)
@@ B.2 Part II (curves)
$\tilde{m}:=\frac{L}{L-A^{(4)}}\left\{\begin{array}[]{llcl}>1&\text{ (meaning $0<1<\tilde{m}$) }&,&L>A^{(4)}\\ \in(0,1)&\text{ (meaning $0<\tilde{m}<1$) }&,&L<0\\ <0&\text{ (meaning $\tilde{m}<0<1$) }&,&0<L<A^{(4)}\end{array}\right.,$ (B.8)
@@
$\tilde{m}:=\frac{R}{R+A^{(4)}}\left\{\begin{array}[]{llcl}>1&\text{ (meaning $0<1<\tilde{m}$) }&,&R<-A^{(4)}\\ \in(0,1)&\text{ (meaning $0<\tilde{m}<1$) }&,&R>0\\ <0&\text{ (meaning $\tilde{m}<0<1$) }&,&-A^{(4)}<R<0\end{array}\right.,$ (B.13)
@@
$\dot{m}=-\frac{\left(R+2A^{(4)}\right)^{2}}{4A^{(4)}}m\left(m-\tilde{m}\right)^{2},$ (B.14)
@@
$\tilde{m}:=\frac{R}{R+2A^{(4)}}\left\{\begin{array}[]{llcl}>1&\text{ (meaning $0<1<\tilde{m}$) }&,&R<-2A^{(4)}\\ \in(0,1)&\text{ (meaning $0<\tilde{m}<1$) }&,&R>0\\ <0&\text{ (meaning $\tilde{m}<0<1$) }&,&-2A^{(4)}<R<0\end{array}\right.,$ (B.18)
the flux diagram is shown in 7.
@@
B.3 Part III (planar regions)
@@
$\dot{m}=-m\left[L-\left(R+2L\right)m+\left(R+L+A^{(4)}\right)m^{2}\right],$ (B.20)
where the expression inside the square brackets in (B.20) is strictly positive. As stated previously, and clear from (B.20), there is a single stationary point, which is $m^{\ast}=0$, and the flux diagram is shown in figure 9.
$<$001111
@@
$m_{+}^{(0)}=\frac{2L}{2L+\left(R-\sqrt{R^{2}-4LA^{(4)}}\right)}<1.$ (B.21)
Hence, $10<m_{-}^{(0)}<m_{+}^{(0)}<$ 1 in region (I).
III.2) Region (II): $R^{2}-4LA^{(4)}>0$, $R+2L>\sqrt{R^{2}-4LA^{(4)}}$ and $R<0$
@@
$m_{-}^{(0)}=\frac{R+2L-\sqrt{R^{2}-4LA^{(4)}}}{2\left(R+L+A^{(4)}\right)}=\frac{2L}{2L-\left(-R-\sqrt{R^{2}-4LA^{(4)}}\right)},$ (B.22)
it follows that $m_{-}^{(0)}>0$ (from the second member of (B.22)); therefore, it is straightforward that $m_{-}^{(0)}>1$ from the last member of (B.22) (note that $-R-\sqrt{R^{2}-4LA^{(4)}}>0$ but the denominator of the last member is positive because $m_{-}^{(0)}>0$). Hence, $0<1<m_{-}^{(0)}<m_{+}^{(0)}$ in region (II).
III.3) Region (III): $R^{2}-4LA^{(4)}>0$, $R+2L<\sqrt{R^{2}-4LA^{(4)}}$, $R+L+A^{(0)}>0$ and $L<0$
@@
$m_{+}^{(0)}=\frac{(-2L)}{(-2L)+\left(\sqrt{R^{2}-4LA^{(4)}}-R\right)},$ (B.23)
since $\sqrt{R^{2}-4LA^{(4)}}-R>0$ for $L<0$, then $10<m_{+}^{(0)}<$ 1. Furthermore,
$m_{-}^{(0)}=\frac{R+2L-\sqrt{R^{2}-4LA^{(4)}}}{2\left(R+L+A^{(4)}\right)}<0,$ (B.24)
@@
$m_{+}^{(0)}=\frac{(-2L)}{(-2L)+\left(\sqrt{R^{2}-4LA^{(4)}}-R\right)}\in(0,1),$ (B.26)
@@
$m_{-}^{(0)}=\frac{(-2L)}{(-2L)-\left(\sqrt{R^{2}-4LA^{(4)}}+R\right)},$ (B.27)
@@
$m_{+}^{(0)}=\frac{2L}{R+2L-\sqrt{R^{2}-4LA^{(4)}}}<0,$ (B.28)
since the numerator and denominator do not share the same sign. Moreover, using the representation (B.19), one sees that $m_{-}^{(0)}>0$ because both the numerator and denominator have the same sign. Finally, from
$m_{-}^{(0)}=\frac{2L}{2L-\left(-R-\sqrt{R^{2}-4LA^{(4)}}\right)},$ (B.29)
one sees that the denominator is positive, because of $m_{-}^{(0)}>0$ and the positivity of the numerator. The conditions $R+L+A^{(0)}<0$ and $L>0$ implies $R<0$. Therefore, $-R-\sqrt{R^{2}-4LA^{(4)}}>0$ for $R<0$ and $L>0$, which leads to $02L>2L-\left(-R-\sqrt{R^{2}-4LA^{(4)}}\right)>$ 0, and, consequently, to $m_{-}^{(0)}>1$. Hence, $m_{+}^{(0)}<0<1<m_{-}^{(0)}$ in region (VI).
@@
C Appendix - case $\overline{A^{(1)}}>0$ and $A^{(4)}=0$
@@
$\dot{n}=-n\left[R-\left(2R+L\right)n+\left(L+R+\overline{A^{(1)}}\right)n^{2}\right],$ (C.2)
where the right-hand side is similar to $F_{0}(m)$ by exchanging $R\leftrightarrow L$ and $A^{(4)}\leftrightarrow\overline{A^{(1)}}$. Furthermore, the roots of the polynomial inside the square brackets in (C.2) are
$n_{\pm}:=\frac{2R+L\pm\sqrt{L^{2}-4R\overline{A^{(1)}}}}{2\left(R+L+\overline{A^{(1)}}\right)},$ (C.3)
@@
Table 4: Division of the $(L,R)$ plane suitable for the construction of the flux diagram of the case $\overline{A^{(1)}}>0$ and $A^{(4)}=0$; the regions in Part VI are defined in picture 11.
C.1 Part IV (points)
@@
$(0,0)$ $\overline{A^{(1)}}\left(1-m\right)^{3}$ 001111$>$
@@ C.2 Part V (curves)
$R>\overline{A^{(1)}}$ $\tilde{m}^{(1)}$001111>>>>>>
$R=0$ $\big{(}L+\overline{A^{(1)}}\big{)}\big{(}1-m\big{)}^{2}\big{(}\tilde{m}^{(1)}-m\big{)}$ $\frac{\overline{A^{(1)}}}{\overline{L+A^{(1)}}}$ $L<-\overline{A^{(1)}}$ $\tilde{m}^{(1)}$001111>>>>>>
$-\overline{A^{(1)}}<L<0$
@@
$L^{2}-4R\overline{A^{(1)}}=0$ $\frac{\big{(}2\overline{A^{(1)}}+L\big{)}^{2}}{4\overline{A^{(1)}}}\big{(}1-m\big{)}\big{(}\tilde{m}^{(1)}-m\big{)}^{2}$ $\frac{\overline{2A^{(1)}}}{\overline{2A^{(1)}}+L}$ $L<-2\overline{A^{(1)}}$ $\tilde{m}^{(1)}$001111>>>>>>
@@
C.3 Part VI (planar regions)
The regions indicated in part VI are shown in figure 11, and the flux diagram is shown in table 13.
@@ C.3 Part VI (planar regions)
$\dot{m}=\left(1-m\right)\left[\overline{A^{(1)}}-\left(2\overline{A^{(1)}}+L\right)m+\left(R+L+\overline{A^{(1)}}\right)m^{2}\right],$ (C.5)
where the expression inside the square brackets in (C.5) is strictly positive. As stated previously, and clear from (C.5), there is a single stationary point, which is $m^{\ast}=1$, and the flux diagram is shown in figure 12.
$>$001111
@@
D Appendix - case $\overline{A^{(1)}}=0$ and $A^{(4)}=0$
@@
$=m\left(1-m\right)\left[\left(R+L\right)m-L\right].$ (D.1)
@@
Table 7: Division of the $(L,R)$ plane suitable for the construction of the flux diagram of the case $\overline{A^{(1)}}=0$ and $A^{(4)}=0$; the regions in Part II are defined in picture 14.
D.1 Part VII (lines)
@@
$R=0$ $-Lm\left(1-m\right)^{2}$ $L>0$ 001111$<$
$L<0$ 001111$>$
$L=0$ $Rm^{2}\left(1-m\right)$ $R>0$ 001111$>$
$R<0$ 001111$<$
$R+L=0$ $-Lm\left(1-m\right)$ $L>0$ 001111$<$
$L<0$ 001111$>$
@@
D.2 Part VIII (planar regions)
The stability analysis of part VIII consists of analyzing six disjoint regions indicated in figure 14.
@@ D.2 Part VIII (planar regions)
In these regions, a third stationary point (apart from $m^{\ast}=0$ and $m^{\ast}=1$), $\tilde{m}^{(01)}$, given by (D.2), emerges. It is straightforward that
$m^{(01)}=\frac{L}{R+L}=\left\{\begin{array}[]{lcl}>1&,&(R+L>0\textnormal{ and }R<0)\textnormal{ or }(R+L<0\textnormal{ and }R>0)\\ \in(0,1)&,&(R+L>0\textnormal{ and }R,L>0)\textnormal{ or }(R+L<0\textnormal{ and }R,L<0)\\ <0&,&(R+L>0\textnormal{ and }L<0)\textnormal{ or }(R+L<0\textnormal{ and }L>0)\end{array}\right.,$ (D.6)

### v2

@@ A Appendix: case $\overline{A^{(1)}}=0$ and $A^{(4)}>0$
Table 4: Division of the $(L,R)$ plane suitable for the construction of the flux diagram of the case $\overline{A^{(1)}}=0$ and $A^{(4)}>0$; the regions in Part III are defined in picture 9.
A.1 Part I (points)
@@ A.2 Part II (curves)
[…]
@@
[…]
@@
$\dot{m}=-\frac{\left(R+2A^{(4)}\right)^{2}}{4A^{(4)}}m\left(m-\tilde{m}\right)^{2},$ (A.14)
@@
[…]
the flux diagram is shown in 8.
@@
A.3 Part III (planar regions)
@@
$\dot{m}=-m\left[L-\left(R+2L\right)m+\left(R+L+A^{(4)}\right)m^{2}\right],$ (A.20)
where the expression inside the square brackets in (A.20) is strictly positive. As stated previously, and clear from (A.20), there is a single stationary point, which is $m^{\ast}=0$, and the flux diagram is shown in figure 10.
<<011
@@
$m_{+}^{(0)}=\frac{2L}{2L+\left(R-\sqrt{R^{2}-4LA^{(4)}}\right)}<1.$ (A.21)
Hence, $0<m_{-}^{(0)}<m_{+}^{(0)}<1$ in region (I).
III.2) Region (II): $R^{2}-4LA^{(4)}>0$, $R+2L>\sqrt{R^{2}-4LA^{(4)}}$, and $R<0$
@@
$m_{-}^{(0)}=\frac{R+2L-\sqrt{R^{2}-4LA^{(4)}}}{2\left(R+L+A^{(4)}\right)}=\frac{2L}{2L-\left(-R-\sqrt{R^{2}-4LA^{(4)}}\right)},$ (A.22)
it follows that $m_{-}^{(0)}>0$ (from the second member of (A.22)); therefore, it is straightforward that $m_{-}^{(0)}>1$ from the last member of (A.22) (note that $R-\sqrt{R^{2}-4LA^{(4)}}>0$ but the denominator of the last member is positive because $m_{-}^{(0)}>0$). Hence, $0<1<m_{-}^{(0)}<m_{+}^{(0)}$ in region (II).
III.3) Region (III): $R^{2}-4LA^{(4)}>0$, $R+2L<\sqrt{R^{2}-4LA^{(4)}}$, $R+L+A^{(0)}>0$, and $L<0$
@@
$m_{+}^{(0)}=\frac{(-2L)}{(-2L)+\left(\sqrt{R^{2}-4LA^{(4)}}-R\right)},$ (A.23)
since $\sqrt{R^{2}-4LA^{(4)}}-R>0$ for $L<0$, then $0<m_{+}^{(0)}<1$. Furthermore,
$m_{-}^{(0)}=\frac{R+2L-\sqrt{R^{2}-4LA^{(4)}}}{2\left(R+L+A^{(4)}\right)}<0,$ (A.24)
@@
$m_{+}^{(0)}=\frac{(-2L)}{(-2L)+\left(\sqrt{R^{2}-4LA^{(4)}}-R\right)}\in(0,1),$ (A.26)
@@
$m_{-}^{(0)}=\frac{(-2L)}{(-2L)-\left(\sqrt{R^{2}-4LA^{(4)}}+R\right)},$ (A.27)
@@
$m_{+}^{(0)}=\frac{2L}{R+2L-\sqrt{R^{2}-4LA^{(4)}}}<0,$ (A.28)
since the numerator and denominator do not share the same sign. Moreover, using the representation (A.19), one sees that $m_{-}^{(0)}>0$ because both the numerator and denominator have the same sign. Finally, from
$m_{-}^{(0)}=\frac{2L}{2L-\left(-R-\sqrt{R^{2}-4LA^{(4)}}\right)},$ (A.29)
one sees that the denominator is positive, because of $m_{-}^{(0)}>0$ and the positivity of the numerator. The conditions $R+L+A^{(0)}<0$ and $L>0$ implies $R<0$. Therefore, $R-\sqrt{R^{2}-4LA^{(4)}}>0$ for $R<0$ and $L>0$, which leads to $2L>2L-\left(-R-\sqrt{R^{2}-4LA^{(4)}}\right)>0$, and, consequently, to $m_{-}^{(0)}>1$. Hence, $m_{+}^{(0)}<0<1<m_{-}^{(0)}$ in region (VI).
@@
B Appendix: case $\overline{A^{(1)}}>0$ and $A^{(4)}=0$
@@
$\dot{n}=-n\left[R-\left(2R+L\right)n+\left(L+R+\overline{A^{(1)}}\right)n^{2}\right],$ (B.2)
where the right-hand side is similar to $F_{0}(m)$ by exchanging $R\leftrightarrow L$ and $A^{(4)}\leftrightarrow\overline{A^{(1)}}$. Furthermore, the roots of the polynomial inside the square brackets in (B.2) are
$n_{\pm}:=\frac{2R+L\pm\sqrt{L^{2}-4R\overline{A^{(1)}}}}{2\left(R+L+\overline{A^{(1)}}\right)},$ (B.3)
@@
Table 5: Division of the $(L,R)$ plane suitable for the construction of the flux diagram of the case $\overline{A^{(1)}}>0$ and $A^{(4)}=0$; the regions in Part VI are defined in Fig. 12.
B.1 Part IV (points)
@@
$(0,0)$ $\overline{A^{(1)}}\left(1-m\right)^{3}$ 011>>
@@ B.2 Part V (curves)
$R>\overline{A^{(1)}}$ $\tilde{m}^{(1)}$011>>>>
$R=0$ $\big(L+\overline{A^{(1)}}\big)\big(1-m\big)^{2}\big(\tilde{m}^{(1)}-m\big)$ $\frac{\overline{A^{(1)}}}{\overline{L+A^{(1)}}}$ $L<-\overline{A^{(1)}}$ $\tilde{m}^{(1)}$011>>>>
$\overline{A^{(1)}}<L<0$
@@
$L^{2}-4R\overline{A^{(1)}}=0$ $\frac{\big(2\overline{A^{(1)}}+L\big)^{2}}{4\overline{A^{(1)}}}\big(1-m\big)\big(\tilde{m}^{(1)}-m\big)^{2}$ $\frac{\overline{2A^{(1)}}}{\overline{2A^{(1)}}+L}$ $L<-2\overline{A^{(1)}}$ $\tilde{m}^{(1)}$011>>>>
@@
B.3 Part VI (planar regions)
The regions indicated in part VI are shown in Fig. 12.
@@ B.3 Part VI (planar regions)
$\dot{m}=\left(1-m\right)\left[\overline{A^{(1)}}-\left(2\overline{A^{(1)}}+L\right)m+\left(R+L+\overline{A^{(1)}}\right)m^{2}\right],$ (B.5)
where the expression inside the square brackets in (B.5) is strictly positive. As stated previously, and clear from (B.5), there is a single stationary point, which is $m^{\ast}=1$, and the flux diagram is shown in figure 13. The flux diagram of the regions of part VI is shown in Fig. 14.
>>011
@@
C Appendix: case $\overline{A^{(1)}}=0$ and $A^{(4)}=0$
@@
$=m\left(1-m\right)\left[\left(R+L\right)m-L\right].$ (C.1)
@@
Table 8: Division of the $(L,R)$ plane suitable for the construction of the flux diagram of the case $\overline{A^{(1)}}=0$ and $A^{(4)}=0$; the regions are defined in Fig. 15.
C.1 Part VII (lines)
@@
$R=0$ $Lm\left(1-m\right)^{2}$ $L>0$ 011<<
$L<0$ 011>>
$L=0$ $Rm^{2}\left(1-m\right)$ $R>0$ 011>>
$R<0$ 011<<
$R+L=0$ $Lm\left(1-m\right)$ $L>0$ 011<<
$L<0$ 011>>
@@
C.2 Part VIII (planar regions)
The stability analysis of part VIII consists of analyzing six disjoint regions indicated in Fig. 15 (see Table 8).
@@ C.2 Part VIII (planar regions)
In these regions, a third stationary point (apart from $m^{\ast}=0$ and $m^{\ast}=1$), $\tilde{m}^{(01)}$, given by (C.2), emerges. It is straightforward that
[…]